OpenStudy (anonymous):

Find an exact value. tan 15° I really appreciate the help! THANK YOU :)

OpenStudy (campbell_st):

use the difference of 2 angles \[\tan(45 - 30) = \frac{\tan(45) - \tan(30)}{1 + \tan(45)\tan(30)}\] now just substitute the exact values that you know

OpenStudy (anonymous):

I got this wrong... I guess the answer isn't 2 - sqrt3 /4

OpenStudy (anonymous):

hi jdoe I appreciate you taking the time to look at my problem :)

OpenStudy (jdoe0001):

$$\bf tan(45^o - 30^o) = \frac{tan(45) - tan(30)}{1 + tan(45)tan(30)}\\ tan(45^o - 30^o) = \cfrac{1 -\frac{1}{\sqrt{3}} }{1 + (1)\left(\frac{1}{\sqrt{3}}\right)}\\ tan(45^o - 30^o) = \large \cfrac{\frac{\sqrt{3}-1}{\sqrt{3}} }{\frac{\sqrt{3}+1}{\sqrt{3}}}\\ $$

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

that isn't one of the options though...argh I missed five questions on this assignment and I want to know what I did wrong before I retake it...want me to show you what the options were?

OpenStudy (jdoe0001):

well, you still need to simplify there further

OpenStudy (jdoe0001):

recall that => \(\bf \large \cfrac{\frac{a}{b}}{\frac{c}{d}} \implies \frac{a}{b} \times \frac{d}{c}\)

OpenStudy (campbell_st):

have you rationalised the denominator you are looking at \[\frac{\sqrt{3} -1}{\sqrt{3} + 1} \times \frac{\sqrt{3} -1}{\sqrt{3} -1} = \frac{3 - 2\sqrt{3} + 1}{3 - 1}\]

OpenStudy (campbell_st):

so the final solution is \[2 - \sqrt{3}\] is that an option

OpenStudy (anonymous):

yes it is! thank you! simplifying! ah got it!

OpenStudy (campbell_st):

so as @jdoe0001 said \[\frac{\frac{\sqrt{3} -1}{\sqrt{3}}}{\frac{\sqrt{3} +1}{\sqrt{3}}} = \frac{\sqrt{3} -1}{\sqrt{3}} \times \frac{\sqrt{3}}{\sqrt{3}+1}\] then you get what I have above

OpenStudy (anonymous):

I'm still trying to see how you got the 2 though? when you simplfied?